half the difference between two numbers is 2.the sum of the greater number and twice the smaller number is 13. find the numbers
Answer by amoresroy(361)

Let x = the smaller number
y = the greater number
(y - x)/2 = 2
y - x = 4
y = 4 + x (1)
y + 2x = 13 (2)
Since y = 4 +x, eq. (2) becomes
(4 + x) + 2x = 13
3x = 13-4
x = 3
y = 4 + 3
y = 7
The numbers are 3 & 7